OpenStudy (anonymous):

find the derivative of the integral of sqrt(t^4-1) from 1 to x.

OpenStudy (anonymous):

According to the fundamental theorem of calc, \[\frac{d}{dx}\int_c^{g(x)}f(t)~dt=f(g(x))\cdot g'(x)\]

OpenStudy (anonymous):

...where \(c\) is a constant, in this case 1, and \(g(x)=x\).